Determine £- {F}. F(s) = -382. - 23s - 41 (s + 3)2(s +4) Click here to view the table of Laplace transforms. Click here to
Answer #1

41=3x 9 + 12B + 4 41=2741BTA ht TB 7c (St 3 1632 =A(876519) TB6475412) tased _38_235-41 Foz (S13) (514) - (35+255+41) A ($13)FS) = 3 1 웁 욺 ما رومی) wob LE 2 따 Sta 와 세혀 ㅍ St2 와 Z 와 사 Nob 니 t772 smte , LI SI 2. th 세 Af] = den2 ? -11 17 lold 1 1, دوم) f​​​​​​